# SPGCL: Effective Graph Contrastive Learning via SVD-Guided Structural Perturbation

### v1

@@ V-E Visualization of Learned Embeddings
GCN, LDS, Geom-GCN, and DropEdge produce embeddings with blurred class boundaries and overlapping clusters.
SPGCL generates well-separated, compact clusters with clearer margins, indicating that the learned representations are more discriminative and structurally coherent.
The visualization confirms that our method better preserves intra-class similarity and inter-class distinction, aligning with the quantitative performance gains.
V-F Analysis of and
We study two hyper-parameters in SPGCL: the base edge generation ratio and an offset in Eq (7).
Setup.
We conduct a 2D grid search over
and report node classification accuracy for each pair. The corresponding heatmaps for Cora and Citeseer are shown in Fig. 3.
Observations.
We observe a clear hill-like (unimodal) trend over the grid. When Top- is too small, the perturbed view is overly similar to the original graph, resulting in limited view diversity. As increases to a moderate range, performance improves and reaches a peak region. However, overly large Top- introduces many low-quality edges, which may cause graph densification and degrade accuracy.
@@ V-G Summary of Experimental Findings
SPGCL consistently outperforms state-of-the-art GNNs and graph structure learning methods across diverse graph types.
@@
Our structured perturbation is superior to naive random perturbations, validating its design principles.
@@
Ablation studies confirm the importance of both stochastic removal and SVD-guided recovery.
@@
Ablating the global consistency loss leads to consistent performance drops on both homophilous and heterophilous datasets (Table IV), verifying that provides complementary global structural alignment beyond node-level contrastive learning.
5.
SPGCL exhibits stronger robustness under increasing structural noise.
6.
Visualization reveals that SPGCL learns more separable and semantically meaningful embeddings.
7.
Sensitivity analysis of shows a clear peak region (Fig. 3), indicating that moderate perturbation is optimal while under- or over-perturbation degrades performance.
These results collectively affirm the effectiveness of SPGCL in learning robust graph representations under structural noise and heterogeneity.
@@ VI Conclusion
In this paper, we addressed the vulnerability of Graph Neural Networks (GNNs) to structural noise by proposing SPGCL, a robust graph contrastive learning framework via SVD-guided structural perturbation. Unlike random perturbations that are structure-agnostic or purely spectral augmentations that may become dense, SPGCL introduces a principled two-stage strategy: it first applies lightweight stochastic edge removal to ensure view diversity, then employs truncated SVD to recover and enhance topologically important edges. This design not only mitigates the risk of accidentally removing critical connections but also enriches the contrastive view with semantically meaningful edges missing in the original graph, while allowing explicit control over structural discrepancy.
Furthermore, we introduced a contrastive fusion module with a global consistency constraint, which aligns node embeddings across views at both local and global structural levels, ensuring robust and invariant representations. Extensive experiments on ten benchmark datasets demonstrate that SPGCL consistently outperforms state-of-the-art GNNs and graph structure learning methods, particularly on heterophilous graphs where structural noise is more detrimental. Ablation studies confirm the necessity of coupling stochastic and spectral perturbations, and robustness analysis shows SPGCL’s superior resilience under increasing structural corruption.
In summary, SPGCL provides an effective, generalizable framework for learning robust graph representations in noisy environments. Future work may explore adaptive perturbation ratios and extend the paradigm to dynamic or heterogeneous graphs.
VII Biography Section
If you have an EPS/PDF photo (graphicx package needed), extra braces are needed around the contents of the optional argument to biography to prevent the LaTeX parser from getting confused when it sees the complicated includegraphics command within an optional argument. (You can create your own custom macro containing the includegraphics command to make things simpler here.)

### v2

@@ V-F Visualization of Learned Embeddings
Baselines (GCN, LDS, Geom-GCN, DropEdge) produce embeddings with blurred class boundaries and overlapping clusters.
SPGCL generates well-separated, compact clusters with clearer margins.
This visualization confirms that SPGCL learns more discriminative and structurally coherent representations, aligning with the quantitative performance gains.
@@ V-G Summary of Experimental Findings
Superior Performance: SPGCL consistently outperforms state-of-the-art GNNs and graph structure learning methods across diverse graph types, proving its general effectiveness.
@@
Design Validity: The structured perturbation is superior to naive random perturbations, and ablation studies confirm the necessity of both stochastic removal and SVD-guided recovery.
@@
Enhanced Robustness: SPGCL exhibits stronger resilience under increasing structural noise, making it suitable for imperfect real-world graphs.
@@
Beneficial Components: The global consistency loss provides complementary structural alignment, and the perturbation scale can be effectively tuned via and .
These results collectively affirm that SPGCL achieves an excellent balance of simplicity in design and power in performance, offering a robust solution for graph contrastive learning. More importantly, they demonstrate how a principled integration of random and spectral augmentations-enabled by a rigorous perturbation framework-can lead to significant empirical gains across diverse graph types and noise levels.
@@ VI Conclusion
This paper proposes SPGCL, a simple yet powerful framework for robust graph contrastive learning via SVD-guided structural perturbation. SPGCL directly addresses the long-standing dichotomy in graph augmentation strategies: random perturbations are structure-agnostic, while purely spectral methods often produce dense and deterministic views. The key to our solution lies in recognizing that the recently developed generalized singular value perturbation theory [4]-which extends classical structural perturbation [12] from symmetric to arbitrary graphs-can serve as the missing bridge to connect these two disparate paradigms.
Building upon this theoretical foundation, we design a principled two-stage strategy that first applies lightweight stochastic edge removal to ensure view diversity, and then employs truncated SVD with perturbation-guided scoring to recover and enhance topologically important edges. This concise yet theoretically grounded design delivers three key advantages: (1) robustness to accidental edge deletion, (2) enrichment with missing high-order connections, and (3) explicit control over structural discrepancy. Furthermore, we introduce a contrastive fusion module with a global consistency constraint to align node embeddings across views at both local and global levels, ensuring invariant and structurally coherent representations.
Extensive experiments on ten benchmark datasets demonstrate that SPGCL consistently outperforms state-of-the-art GNNs and graph structure learning methods. It achieves particularly significant gains on challenging heterophilous graphs while maintaining strong performance on homophilous ones. Ablation studies confirm the necessity of both stochastic and spectral components, and robustness analyses show SPGCL’s superior stability under increasing structural noise. The visualization of learned embeddings further corroborates that SPGCL produces more discriminative and well-separated clusters.
In summary, SPGCL provides an effective and generalizable solution for learning robust graph representations in noisy and heterogeneous environments. Our work also illustrates the versatility of the generalized SVD perturbation framework, showcasing how a method originally developed for graph structure learning can be creatively adapted to solve a fundamentally different challenge in contrastive learning. Future work may explore several promising directions: (1) adaptive perturbation strategies that can dynamically adjust the augmentation strength (e.g., drop ratio , recovery offset ) based on graph properties (e.g., homophily level) or training dynamics, moving beyond static hyper-parameter tuning; (2) extending the paradigm to dynamic graphs (where structure evolves over time) and heterogeneous graphs (with multiple node/edge types), which requires re-thinking structural perturbations in time-aware and relation-aware manners; and (3) investigating the integration of SPGCL with other advanced self-supervised learning paradigms for even broader applicability.
Hao Deng received the B.S. degree in electronic and information engineering from Southwest Petroleum University, Chengdu, China, in 2024. He is currently pursuing the M.S. degree in control science and engineering with Xidian University, Xi’an, China. His current research interests include graph neural networks and graph structure learning. |
Zhang Guo (Member, IEEE) received the B.S., M.S., and Ph.D. degrees in electrical engineering from Xi’an Jiaotong University, Xi’an, China, in 2013, 2016 and 2021, respectively. He is a Member of the IEEE. He is currently an Assistant Professor at the School of Artificial Intelligence, Xidian University. His main research interests include spiking neural network, neuromorphic computing and artificial intelligence. |
Shuiping Gou (Senior Member, IEEE) received the B.S. and M.S. degrees in computer science and technology and the Ph.D. degree in pattern recognition and intelligent systems from Xidian University, Xi’an, China, in 2000, 2003, and 2008, respectively. She is currently a Professor with the Key Laboratory of Intelligent Perception and Image Understanding of Ministry of Education of China, Xidian University. Her research interests include machine learning, data mining, and medical image analysis. |
Bo Liu (Member, IEEE) received the B.A. degree in biotechnology from Zhejiang University, Hangzhou, China, in 2001, and the Ph.D. degree in applied mathematics from Fudan University, Shanghai, China, in 2010. He was a Post-Doctoral Fellow with School of Life Sciences, Fudan University from 2010 to 2012, and a Post-Doctoral Researcher with the Institute of Industrial Science, The University of Tokyo from 2012 to 2014. Since 2014, he has been a professor with Key Laboratory of Intelligent Perception and Image Understanding of the Ministry of Education of China, School of Electronic Engineering (School of Artificial Intelligence since November, 2017), Xidian University, Xi’an. His research interests are the theory and applications of networked systems, including dynamical complex networks, neural networks and deep learning. |